OpenStudy (anonymous):

An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air.

OpenStudy (anonymous):

Use s = vt (s is same in both cases so equate for v)

OpenStudy (ash2326):

Let's the speed of plane be x and wind speed is given as 40 mph Against wind it's given that plane too 3 hours for the trip Return trip with the wind too 2 hours. Both the cases distance covered is the same We know \[distance={speed} \times {time}\] Speed against the wind of the plane is x-40 Speed with the wind is x+40

OpenStudy (ash2326):

So \[ (x-40) \times 3 = (x+40) \times 2\] Do you understand this? Can you solve now?

OpenStudy (anonymous):

so i plug in the speed i think it is for x.. and its supposed to equal eachother?

OpenStudy (ash2326):

Both the sides is distance, which is equal for both the trips

OpenStudy (anonymous):

so its 200! ((200)−40)×3=((200)+40)×2 480=480???? is that right?

OpenStudy (ash2326):

Yeah:D It's right. Speed of plane is 200mph

OpenStudy (anonymous):

thank you! :D im posting another question do you think you can help me?